# Slot-Guided Adaptation of Pre-trained Diffusion Models for Object-Centric Learning and Compositional Generation

### v2

@@ 2.1 Unsupervised Object-Centric Learning
Unsupervised object-centric learning aims to decompose visual scenes into meaningful object representations without need for annotation. Early methods used iterative inference and CNN decoders to reconstruct scenes from object-specific feature vectors (slots) (Eslami et al., 2016; Burgess et al., 2019; Greff et al., 2019; Lin et al., 2019; Jiang et al., 2019; Lin et al., 2020). Attend-Infer-Repeat (AIR) (Eslami et al., 2016) and Sequential AIR (SQAIR) (Kosiorek et al., 2018) reconstructed objects in canonical poses using patch-based decoders. Slot attention (Locatello et al., 2020) and SAVi (Kipf et al., 2021) employed the spatial broadcast decoder (Watters et al., 2019) to predict images and segmentation masks from slots, combined via alpha compositing. While effective on synthetic datasets like CLEVR (Johnson et al., 2017) and 3D Shapes (Burgess & Kim, 2018), these methods struggle with complex, real-world images. Transformer-based decoders, such as SLATE (Singh et al., 2021) and STEVE (Singh et al., 2022), pre-train a discrete VAE (dVAE) to tokenize images and use slot-conditioned transformers for autoregressive reconstruction. Despite improvements, they face challenges in the high-quality reconstruction of complex real images (Singh et al., 2021; Wu et al., 2023a). Methods like DINOSAUR (Seitzer et al., 2023) bypass reconstruction by using self-supervised learning to discover objects but lack image generation capabilities. A relatively recent research direction in OCL is to employ diffusion models as slot decoders, enhancing scene decomposition and visual fidelity in real-world scenarios  (Jiang et al., 2023; Wu et al., 2023b). Several approaches have tackled compositional scene generation through different mechanisms: combining global and symbolic latent variables (Jiang & Ahn, 2020), using hierarchical VAEs with slot attention (Wang et al., 2023), and leveraging hierarchical discrete representations with autoregressive transformers (Wu et al., 2024). While these works focus on synthetic datasets, our work can achieve compositional generation in real-world scenarios.
@@ 4 Experiments
Baselines: We compare SlotAdapt with unsupervised state-of-the-art object-centric methods. On the MOVi-E dataset, we compare with SLATE (Singh et al., 2021), SLATE+, where SLATE’s low capacity dVAE is replaced by a pre-trained VQGAN model (Esser et al., 2021) and Latent Slot Diffusion (LSD) (Jiang et al., 2023). On real-world datasets, we compare with Slot attention (SA) (Locatello et al., 2020), SLATE (Singh et al., 2021), DINOSAUR (Seitzer et al., 2023), LSD (Jiang et al., 2023) and SlotDiffusion (Wu et al., 2023b). In all our experiments on real-world datasets, we use a frozen DINO (Caron et al., 2021) as the visual encoder for all models.
@@ Appendix A Appendix
  • Comparison with GLASS (section A.4): We compare SlotAdapt with GLASS, a similar concurrent study, to show how our work fits into current research.
@@ A.1 Training and Architectural Details
This section provides comprehensive information on the architectural components and training procedures of our model. Moreover, we include the code and training scripts in the supplementary material as a reference.
@@ A.1.1 Architectural Details
We will release our code upon publication and have included it in the supplementary materials for reference.
@@ A.2 Additional Experiments
Synthetic Dataset Results: We evaluate the object discovery and segmentation performance of our method on MOVi-E in comparison to the baseline methods in Table 5, which is presented in Appendix, based on the FG-ARI, instance-level mBO and instance-level mIoU metrics. To measure the representational quality of the learned slots for downstream tasks, we implement a straightforward approach using a 2-layer MLP to predict discrete categories, object positions and 3D bounding boxes. We also display the resulting prediction accuracies in Table 5. We observe significant improvements in almost all metrics, particularly with an approximate 10% enhancement in object discovery and segmentation accuracy.
@@
Table 6: Comparison of segmentation methods on real-world data.
@@
Slot Number: We investigated the impact of slot count by comparing our original configuration of 7 slots with an increased count of 80 slots, departing from the convention established in previous works Seitzer et al. (2023); Wu et al. (2023b); Jiang et al. (2023) where slot count typically matches the maximum number of objects per scene in a given dataset. Our experiments revealed that dramatically increasing the slot count to 80 led to degraded performance, presented in Table 8.
Table 7: Comparison of different slot numbers.
Table 8: Comparison of the number of slots in Slot Attention.
@@
Table 9: Comparison of different starting points of guidance losses.
@@
Additional Token for Slot Attention as a register token: We investigated an alternative approach to capturing global scene information in our architecture. Instead of using the average of all slot tokens as a global register token, we modified the slot attention module to include an additional dedicated slot specifically designed to capture global scene context. This dedicated slot was then directly used in the cross-attention layer. Our experimental results revealed interesting trade-offs: while the dedicated global slot approach improved foreground segmentation quality, it showed slightly decreased performance in semantic segmentation and generation quality. We hypothesize that this performance difference stems from the competitive nature of attention in the slot attention module—the dedicated global slot must compete with object slots for attention weights, potentially limiting its capacity to capture comprehensive scene information compared to our original approach of averaging all slot representations. This suggests that using the collective information from all slots through averaging provides a more robust global representation than a single learned global token, presented in Table 10
Table 10: Comparison of different global information capturing strategies.
@@
Table 11: Effect of CFG values on generation quality metrics.
@@ A.3 Additional Qualitative Examples
Figure 9: Generation Results. We show sample images reconstructed by SlotAdapt conditioned on slots on COCO.
Figure 10: Generation Results. We show sample images reconstructed by SlotAdapt conditioned on slots on VOC.
Figure 11: Segmentation comparisons with other methods We show visualizations of predicted segments on COCO dataset. Compared to other models, our model predicts more complete masks without segmenting objects into parts.
@@
Figure 13: Segmentation comparisons with other methods. We show visualizations of predicted segments on VOC dataset. Compared to other models, our model predicts more complete masks without segmenting objects into parts.
@@
Figure 15: Qualitative comparison: with vs. without guidance. We visualize generated images and predicted segments on COCO dataset. Guidance helps to generate better aligned objects and to solve the “part-whole” hierarchy problem in object segmentation task.
Figure 16: Compositional Generations and Editing. We show visualizations of generated images when the slots are manipulated. In the first two rows, we exchange the highlighted slots. In the middle row, we simply add the highlighted slot to the other image. In the last two rows, we delete the highlighted slots to manipulate the image. For all scenarios, our model can successfully generate realistic images in complex scenarios. This shows our model’s ability to compositional generation and editing.
A.4 Comparison with GLASS
This section presents a comprehensive comparison between our method and the concurrent work GLASS (Singh et al., 2024), highlighting both quantitative and qualitative differences.
Methodological Distinctions: GLASS leverages extra information, such as class labels or image captions, to enhance its performance. While this approach yields certain advantages, it introduces limitations. Primarily, GLASS struggles to differentiate between instances of the same class due to its reliance on semantic masks as pseudo ground truth. Additionally, the need for additional information restricts the applicability of GLASS in fully unsupervised scenarios. In contrast, our method operates without any external supervision, successfully segments individual instances even within the same class, and captures nuanced scene information without relying on pre-defined semantic categories.
@@ A.4 Comparison with GLASS
Table 12 presents a quantitative comparison of our method with GLASS. In terms of semantic overlap, GLASS excels, which is attributable to its use of semantic masks for supervision. However, on complex datasets like COCO, which features multiple instances per image, SlotAdapt achieves comparable performance without any additional supervision. These results underscore the robustness and versatility of our unsupervised approach, particularly in handling complex, multi-instance scenarios. Our method demonstrates that high-quality object segmentation can be achieved without relying on external supervision, offering a more flexible and generalizable solution for diverse image understanding tasks.
Table 12: Unsupervised object segmentation comparisons with the concurrent work GLASS on VOC (left) and COCO (right). We would like to point out that both GLASS and GLASS use extra supervision such as class labels or image caption.

### v3

@@ 2.1 Unsupervised Object-Centric Learning
Unsupervised object-centric learning aims to decompose visual scenes into meaningful object representations without need for annotation. Early methods used iterative inference and CNN decoders to reconstruct scenes from object-specific feature vectors (slots) (Eslami et al., 2016; Burgess et al., 2019; Greff et al., 2019; Lin et al., 2019; Jiang et al., 2019; Lin et al., 2020). Attend-Infer-Repeat (AIR) (Eslami et al., 2016) and Sequential AIR (SQAIR) (Kosiorek et al., 2018) reconstructed objects in canonical poses using patch-based decoders. Slot attention (Locatello et al., 2020) and SAVi (Kipf et al., 2021) employed the spatial broadcast decoder (Watters et al., 2019) to predict images and segmentation masks from slots, combined via alpha compositing. While effective on synthetic datasets like CLEVR (Johnson et al., 2017) and 3D Shapes (Burgess & Kim, 2018), these methods struggle with complex, real-world images. Transformer-based decoders, such as SLATE (Singh et al., 2021) and STEVE (Singh et al., 2022), pre-train a discrete VAE (dVAE) to tokenize images and use slot-conditioned transformers for autoregressive reconstruction. Despite improvements, they face challenges in the high-quality reconstruction of complex real images (Singh et al., 2021; Wu et al., 2023a). Methods like DINOSAUR (Seitzer et al., 2023) bypass reconstruction by using self-supervised learning to discover objects but lack image generation capabilities. To improve slot-based autoencoders, Kakogeorgiou et al. (2024) propose self-training and patch-order permutation strategies, enhancing segmentation in complex real-world images. A relatively recent research direction in OCL is to employ diffusion models as slot decoders, enhancing scene decomposition and visual fidelity in real-world scenarios  (Jiang et al., 2023; Wu et al., 2023b). Several approaches have tackled compositional scene generation through different mechanisms: combining global and symbolic latent variables (Jiang & Ahn, 2020), using hierarchical VAEs with slot attention (Wang et al., 2023), and leveraging hierarchical discrete representations with autoregressive transformers (Wu et al., 2024). While these works focus on synthetic datasets, our work can achieve compositional generation in real-world scenarios.
@@ 4 Experiments
Baselines: We compare SlotAdapt with unsupervised state-of-the-art object-centric methods. On the MOVi-E dataset, we compare with SLATE (Singh et al., 2021), SLATE+, where SLATE’s low capacity dVAE is replaced by a pre-trained VQGAN model (Esser et al., 2021) and Latent Slot Diffusion (LSD) (Jiang et al., 2023). On real-world datasets, we compare with Slot attention (SA) (Locatello et al., 2020), SLATE (Singh et al., 2021), DINOSAUR (Seitzer et al., 2023), LSD (Jiang et al., 2023) and SlotDiffusion (Wu et al., 2023b). In all our experiments on real-world datasets, we use a frozen DINOv2 (Caron et al., 2021; Oquab et al., 2023) as the visual encoder for all models.
@@ Appendix A Appendix
  • Comparison with GLASS and SPOT (section A.4): We compare SlotAdapt with GLASS and SPOT, similar studies, to show how our work fits into current research.
@@ A.1 Training and Architectural Details
This section provides comprehensive information on the architectural components and training procedures of our model.
@@ A.2 Additional Experiments
Synthetic Dataset Results: We evaluate the object discovery and segmentation performance of our method on MOVi-E in comparison to the baseline methods in Table 5 based on the FG-ARI, instance-level mBO and instance-level mIoU metrics. To measure the representational quality of the learned slots for downstream tasks, we implement a straightforward approach using a 2-layer MLP to predict discrete categories, object positions and 3D bounding boxes. We also display the resulting prediction accuracies in Table 5. We observe significant improvements in almost all metrics, particularly with an approximate 10% enhancement in object discovery and segmentation accuracy.
@@
Table 6: Impact of a Better Segmentation Method: We evaluate the effect of a better segmentation model on performance by replacing Slot Attention with BOQ-SA, an improved version. Results are presented on the COCO dataset.
@@
Slot Number: We investigate the impact of slot count by comparing our original configuration of 7 slots with an increased count of 80 slots, departing from the convention established in previous works Seitzer et al. (2023); Wu et al. (2023b); Jiang et al. (2023) where slot count typically matches the maximum number of objects per scene in a given dataset. Our experiments revealed that dramatically increasing the slot count to 80 led to degraded performance, presented in Table 7.
Table 7: Impact of Slot Count on Performance: We analyze the effect of varying the number of slots on the COCO dataset.
@@
Table 8: Impact of Guidance Loss Timing: We compare the effect of applying the guidance loss from the start of training versus introducing it after the model has learned initial representations. Results are presented on the COCO dataset.
@@
Additional Token for Slot Attention as a register token: We investigate an alternative approach to capturing global scene information in our architecture. Instead of using the average of all slot tokens as a global register token, we modified the slot attention module to include an additional dedicated slot specifically designed to capture global scene context. This dedicated slot was then directly used in the cross-attention layer. Our experimental results revealed interesting trade-offs: while the dedicated global slot approach improved foreground segmentation quality, it showed slightly decreased performance in semantic segmentation and generation quality. We hypothesize that this performance difference stems from the competitive nature of attention in the slot attention module—the dedicated global slot must compete with object slots for attention weights, potentially limiting its capacity to capture comprehensive scene information compared to our original approach of averaging all slot representations. This suggests that using the collective information from all slots through averaging provides a more robust global representation than a single learned global token, presented in Table 9
Table 9: Impact of Global Information Capturing Strategies: We compare slot averaging and the use of an additional slot token for capturing global information. Results are presented on the COCO dataset.
@@
Table 10: Impact of CFG Value on Generation Quality: We evaluate the effect of different CFG values on generation quality using the COCO dataset.
@@ A.3 Additional Qualitative Examples
Figure 9: Generation Results. We show sample images from COCO, reconstructed by SlotAdapt conditioned on slots.
Figure 10: Generation Results. We show sample images from VOC, reconstructed by SlotAdapt conditioned on slots.
Figure 11: Segmentation comparisons with other methods We show visualizations of predicted segments on COCO dataset. Compared to other models, our model tends to produce more coherent masks with fewer fragmented segments.
@@
Figure 13: Segmentation comparisons with other methods. We show visualizations of predicted segments on VOC dataset. Compared to other models, our model tends to produce more coherent masks with fewer fragmented segments.
@@
Figure 15: Qualitative comparison: with vs. without guidance. We visualize generated images and predicted segments on COCO dataset. Guidance helps to generate better aligned objects and to mitigate the “part-whole” hierarchy problem in object segmentation task.
Figure 16: Compositional Generations and Editing. We show visualizations of generated images when the slots are manipulated. In the first two rows, we exchange the highlighted slots. In the middle row, we simply add the highlighted slot to the other image. In the last two rows, we delete the highlighted slots to manipulate the image. For all scenarios, our model can successfully generate realistic images in complex scenarios. This shows our model’s ability for compositional generation and editing.
A.4 Comparison with GLASS and SPOT
This section presents a comprehensive comparison between our method and the concurrent work GLASS (Singh et al., 2024) and SPOT (Kakogeorgiou et al., 2024), highlighting both quantitative and qualitative differences.
Methodological Distinctions: GLASS leverages extra information, such as class labels or image captions, to enhance its performance. While this approach yields certain advantages, it introduces limitations. Primarily, GLASS struggles to differentiate between instances of the same class due to its reliance on semantic masks as pseudo ground truth. Additionally, the need for additional information restricts the applicability of GLASS in fully unsupervised scenarios. In contrast, our method operates without any external supervision, successfully segments individual instances even within the same class, and captures nuanced scene information without relying on pre-defined semantic categories. We should also note that GLASS does not present any compositional editing results.
SPOT operates in a fully unsupervised manner by introducing (i) an attention-based self-training mechanism that distills improved slot-based attention masks from the decoder to the encoder, and (ii) a patch-order permutation strategy for autoregressive transformers to better utilize slot representations during reconstruction. Although SPOT achieves strong object segmentation performance, particularly on complex real-world images, it still faces challenges in accurately differentiating fine-grained object instances within the same class. We should also note that SPOT does not present any compositional editing or reconstruction results.
@@ A.4 Comparison with GLASS and SPOT
Table 11 presents a quantitative comparison of our method with GLASS and SPOT. In terms of semantic overlap, GLASS excels, which is attributable to its use of semantic masks for supervision. However, on complex datasets like COCO, which features multiple instances per image, SlotAdapt achieves comparable performance without any additional supervision. SPOT, while fully unsupervised, focuses on enhancing slot representations through self-training and patch-order permutation strategies. Although it performs well in capturing global scene structures, particularly on COCO, it lags behind our method in fine-grained instance differentiation, as reflected in the FG-ARI scores.
These results underscore the robustness and versatility of our unsupervised approach, particularly in handling complex, multi-instance scenarios. Our method demonstrates that high-quality object segmentation can be achieved without relying on external supervision, offering a more flexible and generalizable solution for diverse image understanding tasks.
Table 11: Unsupervised object segmentation comparisons with the concurrent work GLASS and SPOT on VOC (left) and COCO (right). We would like to point out that both GLASS and GLASS use extra supervision such as class labels or image caption. SPOT is a fully unsupervised method.